A money box contains one-rupee and two-rupee coins in the ratio 5:6. If the total value of the coins in the money box is Rs. 85, find the number of two-rupee coins.
Answer by TimothyLamb(4379)

p + 2q = 85
p/q = 5/6
p = (5/6)q
(5/6)q + (12/6)q = 85
(17/6)q = 85
q = (6/17)85
q = 30
Answer:
number of two-rupee coins in box is 30
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php